OpenStudy (anonymous):

If (5x-y) varies directly as (2x+3y) which is always a constant?

OpenStudy (anonymous):

whst are the options

OpenStudy (anonymous):

A) x+y, B) x-y, C) xy, D)x/y

OpenStudy (phi):

(5x-y) varies directly as (2x+3y) means there is a constant, we will call it k, so that 5x-y= k(2x+3y) if you distribute the k on the right side, you get 5x -y =2kx +3k y collect terms 5x-2kx = 3k y + y factor out x on the left side, and factor you y on the right side (5-2k) x = (3k+1) y

OpenStudy (phi):

notice that in (5-2k) x = (3k+1) y that 5-2k is a constant number. in other words, k is a fixed number and 5-2k will also be some number that does not change. same for 3k+1 if we divide both sides by (5-2k) we get x= (3k+1)/(5-2k) y divide both sides by y \[ \frac{x}{y}= \frac{3k+1}{5-2k} \]

OpenStudy (anonymous):

Thanks